OpenStudy (perii224):

PLEASE HELP WILL MEDAL The sum of two numbers is 66. The larger number is 2 more than 3 times the smaller. What are the two numbers? 22 and 44 20 and 46 18 and 48 16 and 50 A bag has 6 red marbles, 8 blue marbles and 4 yellow marbles. What is the probability of picking a red marble, replacing it, and then picking a blue marble 1/18 2/9 3/22 4/27

OpenStudy (perii224):

@Mehek14 @dumbcow

OpenStudy (dumbcow):

1) x + y = 66 x = 2+3y ----> (2+3y) + y = 66 solve for y

OpenStudy (perii224):

hmm, im not sure

OpenStudy (dumbcow):

2) Total marbles = 18 prob red = 6/18 = 1/3 prob blue = 8/18 = 4/9 multiply the probabilities -----> (1/3)(4/9) = 4/27

OpenStudy (perii224):

Ty!